OpenStudy (anonymous):

Will medal and fan! Amanda's first four test scores are 77, 95, 84, and 69. What does Amanda need to score on the next test to have at least an 82 average? A) 85 or greater B) 90 or greater C) greater than 85 D) greater than 92

OpenStudy (anonymous):

Thanks for your help in advance! :)

OpenStudy (anonymous):

hmm

OpenStudy (fortytherapper):

Let's find the average of his first 4 test scores. What would that be?

OpenStudy (anonymous):

hmm

OpenStudy (landon34):

do you know how to find the mean?

OpenStudy (anonymous):

yeah add the numbers then divide the number of numbers right?

OpenStudy (landon34):

yeah.

OpenStudy (anonymous):

to the sum of all the numbers?

OpenStudy (landon34):

you add up all the numbers and then divide by the number of numbers

OpenStudy (anonymous):

you add all the scores up and divided them by how many numbers you added up. In this case, you'd add 77, 75, 84, and 69. Then you divided by four because there are four numbers.

OpenStudy (anonymous):

so in this case............. all combined would be 325 divided by 4 would be..............

OpenStudy (anonymous):

total is 81.25

OpenStudy (landon34):

yep.

OpenStudy (anonymous):

81.25!

OpenStudy (landon34):

you got it.

OpenStudy (anonymous):

So what is it asking in this case?

OpenStudy (fortytherapper):

Alright, so you got 81.25. Now if those 5 test are 100% of his grade, what % would each test be worth?

OpenStudy (anonymous):

or what is the answer lol.

OpenStudy (anonymous):

so dividing 100 and 5 would be..............

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

20%, right

OpenStudy (anonymous):

awesomesauce!

OpenStudy (landon34):

yeah

OpenStudy (fortytherapper):

Right! Which would also be .20 out of 1. See how I got that part?

OpenStudy (anonymous):

20 out of 1?

OpenStudy (fortytherapper):

.20, not 20

OpenStudy (anonymous):

.20 ok

OpenStudy (anonymous):

20% = 0.20

OpenStudy (fortytherapper):

So since you got those two numbers, multiply 81.25 and .20 for me That will give you his current grade

OpenStudy (anonymous):

ok because 20% is some of it not all so it would be behind the decimal ok i get it.

OpenStudy (anonymous):

ok.

OpenStudy (fortytherapper):

And right

OpenStudy (anonymous):

16.25?.....

OpenStudy (fortytherapper):

Wait, my fault I meant multiply 81.25 and .80

OpenStudy (anonymous):

oh ok lol

OpenStudy (fortytherapper):

Because he's completed 4/5 test and 4/5 = .80 My bad

OpenStudy (anonymous):

ooh 65%

OpenStudy (fortytherapper):

Yes, so thats what he has

OpenStudy (anonymous):

that is not a good or passing grade XD

OpenStudy (landon34):

lol

OpenStudy (anonymous):

so what is the answer @FortyTheRapper?

OpenStudy (landon34):

in my school, 72% is passing.

OpenStudy (anonymous):

cause 65% not a answer choice @FortyTheRapper

OpenStudy (fortytherapper):

Now the grade he wants is a 82, but he has a 65.. so lets subtract 82 and 65

OpenStudy (anonymous):

17%

OpenStudy (fortytherapper):

That will get him the number of points out of the 20% he needs And right, 17/20

OpenStudy (fortytherapper):

So last step is to convert 20 times ? = 100

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

this is kind of a complicated way of doing it... If the student needs 82% average for 5 tests, each worth 100 points, and previous test scores are 77, 95, 84, and 69: \[\frac{\text{points earned}}{\text{points possible}} \ge 82\%\] But points earned will be \[77+95+84+69 +x\]where \(x\) is the score on the upcoming test. \[\frac{77+95+84+69 +x}{100+100+100+100+100} = 0.82\]will give the lowest score possible to get an \(82\%\) average.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

20 times 5 is 100

OpenStudy (fortytherapper):

Right, so to convert you would do the same to the top 17 times 5 = ? ----- 20 times 5 = 100%

OpenStudy (anonymous):

17 times 5 is 85.

OpenStudy (fortytherapper):

Right, so he would need a 85/100 to get that 82 average

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

so A?

OpenStudy (fortytherapper):

Yep

OpenStudy (anonymous):

ok tysm @FortyTheRapper here is your medal and fan!

OpenStudy (anonymous):

closing question have a good day!

OpenStudy (fortytherapper):

You too